A current of 3.18 A is passed through a Sn ( NO 3 ) 2 solution for 1.20 h . How much tin is plated out of the solution??
Answer #1

Answer:

  •   Equivalent mass of Sn = 118.7/2
  • = 59.35 g
  • Q = it
  • = 3.68*t
  • Therefore
  • No. of Equivalent Deposit = (3.18*1.2)*60*60/96500
  • i/59.35 = (3.18*1.2)*60*60/96500
  • m = 59.35*(3.18*1.2)*60*60/96500
  • = 8.448gm

                                                               Please Rate My Answer.......Thank......u...